# RocqStar: Leveraging Similarity-driven Retrieval and Agentic Systems for Rocq generation

### v1

@@ 5 Related Work
Many Rocq generation methods improve generation using Retrieval Augmentation. Most of those works solve the hint selection problem, described in § 2. Those approaches build proofs tactic by tactic, retrieving relevant lemmas or definitions to use in the next step. The problem of searching for existing proofs that could advance the generation is barely described in the literature. CoqPilot [14] and Rango [27] pack the context for the generator model with theorems most similar to the one we are solving. Our work proposes a novel method of doing premise selection and shows improvement over the baseline from previous works [14, 27].
@@ 6 Conclusion
Our work pioneers the use of Agentic Systems applied to Formal Verification. We have implemented an advanced pipeline that includes rigorous planning via multi-agent debate, domain-specific tooling, and an adaptive executor–critic loop that iteratively refines proofs based on partial progress. We conclude that our RocqStar agent shows promising results, surpassing strong baselines and highlighting the applicability of agentic systems in the domain of theorem proving. We evaluated the idea of applying multi-agent debate to the planning stage of the agentic pipeline in our ablation study and demonstrated that it substantially improves the downstream proof success rate.
@@
We thank Ekaterina Verbitskaia, Ivan Kabashnyi, Maksim Rozenberg, and Pavel Guliaev for their valuable comments regarding this work.
@@ Appendix A IMM Evaluation Dataset
The collected IMM-300 dataset from the CoqPilot [14] work includes only thereoms with proofs of length no more than 20. For that reason the bucket with the most difficult theorems is labeled 9 — 20 tactics. This decision has been made, reflecting CoqPilot’s original focus on subgoals and shorter lemmas. Theorems of length no more than 20 tactics account for 83% of all proofs in the IMM project. As we take the same dataset, it possesses the same limitations. Therefore, we have not evaluated our solution on theorems, for which the reference proof contains more than 20 tactics. However, such theorems are quite rare.
The exact list of theorems used in each group could be found in the repository of the CoqPilot project: https://github.com/JetBrains-Research/coqpilot/blob/main/etc/docs/benchmark/.
@@ Appendix B Encoder Training Dataset
all: try (left; auto) || (right; auto).
@@
The dataset is stored as a collection of JSON files and, due to its relatively small size, is stored within the repository, in the sub-directory with the model training code: https://github.com/JetBrains-Research/rocqstar-rag/tree/main/proof-embeddings/data/.
Dataset is split into training, validation, and test sets with proportions of 70%, 20%, and 10% respectively. Theorems from the same file do not appear in different sets. Parameters of building the dataset are listed in Table 4b. Pair of statements is considered as negative, if the proof_distance between them is greater than 0.65, and positive if it is less than 0.3. If the diatance is in range $[0.45,0.65]$, with probability of 30% it is also considered to be a negative pair (see hard negatives in § 2.2).
@@ Appendix C Encoder Details
algorithm AdamW (0.9, $0.9,\;0.99,\;\mathrm{e}{-2}$)
@@
batch size (stmts) 32
@@
Hyperparameters used for training the embedder model are listed in Table 4. We have used microsoft/codebert-base as the base model and trained our embedder for 22000 steps with a batch size 32. We applied a dropout of 0.1 on the last layer of the model; the embedding dimension is 768, and the maximum sequence length is 128. We use AdamW optimizer with a linear warmup schedule for 10% of the training steps.
@@
If we measure the distance between theorems from Firgure 3 using the convential Jaccard distance, which is used by default in CoqPilot, we get $0.67$:
@@ Appendix D Agent toolset
  • get_theorem_names: Retrieves available theorem names from a file, including the target theorem.
  • get_theorem_names_excl: Retrieves available theorem names from a file with target theorem excluded from the list.
  • get_current_target_state: Returns the stage of the proof for the target theorem in the current session.
@@
  • get_objects: Returns output of Rocq’s Print All command, issued in the context of the current session. This command prints all defined objects in the current file. In particular, that would mean printing all statements of theorems available above the one we are trying to prove at the moment of request.
@@
  • check_term: Checks a term in the current session’s file. Uses Rocq’s Check command. Similar to print_term or about_term, but outputs only the type of the term. In the case of a theorem, it outputs its statement.
@@ Appendix E Plan refinement via Multi-Agent Debate
The proposed proof strategy is quite vague, too wordy, and imprecise. With this plan, the executor agent started by fetching similar theorems and then got stuck in a cycle of ad-hoc attempts to check incorrect proofs. The winning plan, on contra, was precise, concise, and clear. It had a top-level structure of the proof incorporated into the plan and made a winning solution to use the search command, which helped to finish the proof:
@@ Appendix F Experiments compute resources
Embedder training  During our final training run of the embedding model for RocqStar ranker, the model consumed roughly 43 GB of GPU-process memory and only about 6% of the host’s RAM. Over 13.5 hours on a single NVIDIA H100 accelerator (with 20 CPU cores and 160 GB of system memory), disk usage grew steadily from 28 GB to 76 GB as checkpoints and logs accumulated. GPU utilization stabilized above 85% shortly after the warmup phase and remained near saturation for the remainder of training. To sum up, our setup runs comfortably on a single GPU node with modest additional CPU and memory overhead.
Embedder model evaluation  Experiments were conducted on a single MacBook Pro with an M1 chip. The only computationally expensive part of the experiments (for the local machine) is launching multiple Coq-LSP servers at once (CoqPilot benchmark does that to optimize the time of the experiments and accelerate type-checking). As we use a middleware service over LLM APIs, our financial estimations might not be accurate. However, we roughly estimate 12 generation attempts per theorem with seven contextual theorems at 12 cents per theorem for Claude 3.5 and 7 cents for GPT-4o. We run an experiment for 300 theorems and repeat it three times, resulting in 114 US Dollars.
Agent evaluation  In the case of the agent evaluation, we ran the experiment only once and did not provide the confidence intervals due to financial limitations. We ran our agent on the IMM-300 dataset, and afterward, we compared two versions of the agent on the IMM-50 dataset in our § 4.3. That results in 400 attempts to prove different theorems. We estimate a single attempt at 1.3 US dollars. Therefore, we estimate the cost of the evaluation of the agent to be 520 US dollars.

### v2

@@ 5 Related Work
Many Rocq generation methods improve generation using Retrieval Augmentation. Most of those works solve the hint selection problem [3, 28], described in §˜2. Those approaches build proofs tactic by tactic, retrieving relevant lemmas or definitions to use in the next step. The problem of searching for existing proofs that could advance the generation is barely described in the literature. CoqPilot [15] and Rango [28] pack the context for the generator model with theorems most similar to the one we are solving. Our work proposes a novel method of doing premise selection and shows improvement over the baseline from previous works [15, 28].
@@ 6 Conclusion
Our work pioneers the use of Agentic Systems applied to Formal Verification. We have implemented an advanced pipeline that includes rigorous planning via multi-agent debate, domain-specific tooling, and an adaptive executor–critic loop that iteratively refines proofs based on partial progress. We conclude that our RocqStar agent shows promising results, surpassing strong baselines and highlighting the applicability of agentic systems in the domain of theorem proving. The ablation study further demonstrates that both multi-agent planning and reflection are key to maintaining stable reasoning and achieving consistent improvements in the proof’s success rate.
@@
We thank Ekaterina Verbitskaia, Ivan Kabashnyi, Maksim Rozenberg, and Pavel Guliaev for their valuable feedback on this work. We are also grateful to the JetBrains Koog team for providing a hands-on and well-designed framework for building AI agents. Finally, we thank the Dynamic Program Analysis Team (Egor Klimov, Nikita Dukin, and Saga Rut Sunnevudóttir) for their help in analyzing the agent’s execution traces.
@@ Appendix A IMM Evaluation Dataset
The collected IMM-300 dataset from the CoqPilot [15] work includes only theorems with proofs of length no more than 20. For that reason the bucket with the most difficult theorems is labeled 9 — 20 tactics. This decision has been made, reflecting CoqPilot’s original focus on subgoals and shorter lemmas. Theorems of length no more than 20 tactics account for 83% of all proofs in the IMM project. As we take the same dataset, it possesses the same limitations. Therefore, we have not evaluated our solution on theorems, for which the reference proof contains more than 20 tactics. However, such theorems are quite rare.
The exact list of theorems used in each group could be found in the repository of the CoqPilot project: https://github.com/JetBrains-Research/coqpilot/blob/main/etc/docs/benchmark.
@@ Appendix B Encoder Training Dataset
all: try (left; auto) $|$ (right; auto).
@@
The dataset is stored as a collection of JSON files and, due to its relatively small size, is stored within the repository, in the sub-directory with the model training code: https://github.com/JetBrains-Research/rocqstar-rag/tree/main/proof-embeddings/data.
Dataset is split into training, validation, and test sets with proportions of 70%, 20%, and 10% respectively. Theorems from the same file do not appear in different sets. Parameters of building the dataset are listed in Table˜4b. Pair of statements is considered as negative, if the proof_distance between them is greater than 0.65, and positive if it is less than 0.3. If the distance is in range $[0.45,0.65]$, with probability of 30% it is also considered to be a negative pair (see hard negatives in §˜2.2).
@@ Appendix C Encoder Details
algorithm AdamW ($0.9,\;0.99,\;\mathrm{e}{-2}$)
@@
batch size (stmts) 16
@@
The hyperparameters used for training the embedder model are listed in Table˜4. We have used microsoft/codebert-base as the base model and trained our embedder for 22000 steps with a batch size 16, since we use many negative samples in the loss. We applied a dropout of 0.1 on the last layer of the model; the embedding dimension is 768, and the maximum sequence length is 128. We use AdamW optimizer with a linear warmup schedule for 10% of the training steps.
@@
If we measure the distance between theorems from Figure 3 using the conventional Jaccard distance, which is used by default in CoqPilot, we get $0.67$:
@@ Appendix D Agent toolset
  • get_theorem_names: Retrieves theorem names available in the file, including the target theorem.
  • get_theorem_names_excl: Retrieves the available theorem names from a file with the target theorem excluded from the list.
  • get_current_target_state: Returns the proof stage for the target theorem in the current session.
@@
  • check_term: Checks a term in the current session’s file. Uses Rocq’s Check command. Similarly to print_term or about_term, but outputs only the type of the term. In the case of a theorem, it outputs its statement.
@@ Appendix E Plan refinement via Multi-Agent Debate
The proposed proof strategy is quite vague, too wordy, and imprecise. With this plan, the executor agent started by fetching similar theorems and then got stuck in a cycle of ad-hoc attempts to check incorrect proofs. In contrast, the winning plan was precise, concise and clear. It had a top-level structure of the proof incorporated into the plan and made a winning solution to use the search command, which helped to finish the proof:
@@ Appendix F Experiments compute resources
Embedder training  During our final training run of the embedding model for RocqStar ranker, the model consumed roughly 43 GB of GPU-process memory and only about 6% of the host’s RAM. Over 31 hours on a single NVIDIA H200 accelerator (with 16 CPU cores and 200 GB of system memory), disk usage grew steadily from 28 GB to 76 GB as checkpoints and logs accumulated. GPU utilization stabilized above 90% shortly after the warmup phase and remained near saturation for the remainder of training. To sum up, our setup runs comfortably on a single GPU node with modest additional CPU and memory overhead.
Embedder model evaluation  Experiments were conducted on a single MacBook Pro with an M1 chip. The only computationally expensive part of the experiments (for the local machine) is launching multiple Coq-LSP servers at once (CoqPilot benchmark does that to optimize the time of the experiments and accelerate type-checking). As we use a middleware service over LLM APIs, our financial estimations might not be accurate. However, we roughly estimate 12 generation attempts per theorem with seven contextual theorems at 12 cents per theorem for Claude 3.5 and 7 cents for GPT-4o. Running the experiment on 300 theorems and repeating it three times amounts to a total of 171 US Dollars. Performing the same experiment for three different retrieval engines brings the overall cost to approximately 513 US Dollars.
Agent evaluation  In the case of the agent evaluation, we ran the experiment only once and did not provide the confidence intervals due to financial limitations. We ran our agent on the IMM-300 dataset, and afterward, we compared five versions of the agent on the IMM-50 dataset in our §˜4.3. That results in 550 attempts to prove different theorems. We estimate a single attempt at 1.3 US dollars. Therefore, we estimate the cost of the evaluation of the agent to be 715 US dollars.